OpenStudy (anonymous):

Find an equation perpendicular to x=-5 and passing through (7,5)

OpenStudy (anonymous):

As it's perpindicular, the gradient will be reciprocal: 1over 5. Then, use the formula for finding a line equation (y-y1)=m(x-x1)! Substituting your co-ordinates in for x and y, that gives (y-5)=m(x-7). M, as we've seen, is 1 over 5. That gives you (y-5)=1/5(x-7). Assuming you're still with me, multiply it out and you have y=1/5x+(18/5), in the form y-mx+c

OpenStudy (anonymous):

also here you might find something useful :) http://www.okc.cc.ok.us/maustin/EquationsofLine/Equations%20of%20a%20Line.htm

OpenStudy (anonymous):

it's y=mx+c ;)

OpenStudy (anonymous):

Whoops, typo!

OpenStudy (anonymous):



OpenStudy (anonymous):

Oh snap. y=5